OpenStudy (anonymous):

10 mathstina In a methane(CH4). The position of the hydrogen atoms are (0,0,0),(1,1,0),(1,0,1)and (0,1,1);the positionof the carbon atom is (0.5,0.5,0.5). Compute the angle HCH or the angle formed by the hydrogen-carbon-hydrogen bond.

OpenStudy (anonymous):

calculus?

OpenStudy (anonymous):

yes,vectors.

OpenStudy (anonymous):

nope 3d

OpenStudy (anonymous):

yeah 3d and vectors do the same

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

[-0.5,-0.5,-0.5] . [0.5,-0.5,0.5] = 2*sqrt(0.75) cos(A) =>arccos(-.25/(2*sqrt(0.75)) = A

OpenStudy (anonymous):

the angle is 98.3 degrees?

OpenStudy (anonymous):

atchyut is the ans right?

OpenStudy (anonymous):

actually i donno but i found this once take a look of this

OpenStudy (anonymous):

Since we want the angle between ANY H-C-H bond, choose any two vertices of the tetrahedron to be your two hydrogen atoms (the angles are all the same, so it doesn't matter which two you choose). Here, I chose (0, 0, 1) and (0, 1, 0) to be the two hydrogen atoms (they seem to be the easiest to pick). We want the angle between the following two vectors: (a) one with endpoints <1/2, 1/2, 1/2> and <0, 0, 1> (b) one with endpoints <1/2, 1/2, 1/2> and <0, 1, 0>. Subtracting head from tail, the two vectors are: <1/2, 1/2, 1/2 - 1> and <1/2, 1/2 - 1, 1/2> ==> <1/2, 1/2, -1/2> and <1/2, -1/2, 1/2>. The two vectors above both have magnitudes √(1/4 + 1/4 + 1/4) = √3/2. Using: a • b = |a||b|cosθ, and: a • b = <1/2, 1/2, -1/2> • <1/2, -1/2, 1/2> = (1/2)(1/2) + (1/2)(-1/2) + (-1/2)(1/2) = 1/4 - 1/4 - 1/4 = -1/4, we have: θ = arccos[(a • b)/(|a||b|)] = arccos{(-1/4)/[(√3/2)(√3/2)]} ≈ 109.5°, which is to be expected. I hope this helps!

OpenStudy (anonymous):

ok. Thanks. is this the simpler method?

OpenStudy (anonymous):

yeah best method but the points here varies and not the same points given by you

OpenStudy (anonymous):

(1,0,0), (0,1,0), (0,0,1), and (1,1,1). the centroid is at (1/2,1/2,1/2).

OpenStudy (anonymous):

using dot product is always best and easier way than to remember some formulae

OpenStudy (anonymous):

ok. i'll try using using my points .

OpenStudy (anonymous):

Thats good goood luck and do post the answer

OpenStudy (anonymous):

I will check with mine and we can have the correct answer then

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Answer will be the same

OpenStudy (anonymous):

atchyut i got the same ans , using my points. thanks a lot!!!!

OpenStudy (anonymous):

109 degrees itself

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

yes, 109.47 degrees.

OpenStudy (anonymous):

atchyut cani ask u one more qn on vectors?

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

two planes with eqns 3x+4y=1 & x+y-z=3 a) find the angle btwnthe planes. b)determine the symmetric eqns fot the line of intersection L of these two planes.

OpenStudy (anonymous):

i have done part a), i got 36.07 degrees. am i right? hw abt part b)?

OpenStudy (anonymous):

hmmm i am not so sure about it... But symmetric equationis something like (x-x1)/a=(y-y1)/b=(z-z1)/c isnt it???

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

hmmmm then can u find the point of intersection?????

OpenStudy (anonymous):

no, i'm nt sure hw to find?

OpenStudy (anonymous):

Yeah listen in order to find a point in this situation just put z=0 and solve the other two equations then it will be helpful not that this is the exact procedure but this is taught to me by some prof and so it will be helpful sometimes as in this case

OpenStudy (anonymous):

Keeping z=0 and solving other two u will get (11,-8,0) which lies on both the planes

OpenStudy (anonymous):

ok, then?

OpenStudy (anonymous):

It will be helpful because they are planes and u will definitely find a pint where it meet x,y plane and thats z=0 understood

OpenStudy (anonymous):

and now find any two points on each planes get the directional ratios and substitute them in a,b,c

OpenStudy (anonymous):

is the ans: x=11-4t y=-8+3t z=-t

OpenStudy (anonymous):

Yeah right if d.r's are correct

OpenStudy (anonymous):

i used cross product to get the ans. was that ur method?

OpenStudy (anonymous):

not exactly mine but yeah answer matched

OpenStudy (anonymous):

if possible can teach me ur method?

OpenStudy (anonymous):

First tell me about ur cross product method dude coz i am confused how we can do that

OpenStudy (anonymous):

cross product using 2 eqns, get -4i+3j-k. then using (11,-8,0). so,x=11-4t,y=-8+3t,z=-t.

OpenStudy (anonymous):

Yeah u can do it and it is the good procedure... Mine is somewhat trialing type.. For theoretical proofs ur's is correct and mine is for to choose answer from given options. yeah u can follow that and that's right

OpenStudy (anonymous):

ok,thanks for ur time n help.

OpenStudy (anonymous):

when we get that point (11,-8,0) it is (x1,y1,z1) and a b c which are called d.r's are obtained by ur cross product then symmetric equations are x-11/-4 = y+8/3 =z-0/-1 = t and each of the equations u have written for x y z are called parametric equations in t. ie, for each value of t u will get each point on the line which is intersection of the two given planes

OpenStudy (anonymous):

is parametric eqn & symmetric eqn has same meaning? am i right?

OpenStudy (anonymous):

thanks n bye!

OpenStudy (anonymous):

sorry dude sys restart

OpenStudy (anonymous):

Actually no symmetric equation is what i told u in the first (x-x1)/a=(y-y1)/b=(z-z1)/c while parametric eqns are individually for each of x y and z x=11-4t,y=-8+3t,z=-t.

OpenStudy (anonymous):

ok.understood, i learnt a lot

OpenStudy (anonymous):

fine good luck

OpenStudy (anonymous):

which country r u frm?

OpenStudy (anonymous):

BYE

OpenStudy (anonymous):

i am from India

OpenStudy (anonymous):

i'm frm singapore. nice to meet u

OpenStudy (anonymous):

there should be abt 1pm?? here 3:27 pm

OpenStudy (anonymous):

yeah thats right

OpenStudy (anonymous):

and what are u studying in btw?

OpenStudy (anonymous):

in college.r u a student too?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok i'veto go out .bye

OpenStudy (anonymous):

k k tc bubye